OpenStudy (anonymous):

Calculate the work against gravity required to build a structure out of brick assuming brick has a density of 80 lb/ft^3. The structure is a tower of height 20 ft and a square base of side 10 ft

OpenStudy (amistre64):

perimeter, volume, what are we bricking in?

OpenStudy (anonymous):

We are building a pyramid. I am just having issues because I am unsure if I should express the area first as l^2 = (10)^2 = 100 or, if I should use similar triangles to set it as 5/20 = (l/2)/(20-y).

OpenStudy (anonymous):

So then my l^2 would be (1/4)*(20-y)

OpenStudy (amistre64):

so its just the lateral area?

OpenStudy (amistre64):